# Auditing a Dutch Public Sector Risk Profiling Algorithm Using an Unsupervised Bias Detection Tool

### v2

@@ Appendix B Clustering Results for CUB 2019 data
We follow the exact same procedure as outlined in Section 4, only changing the options for the minimum number of samples to $1,000,2,000,3,500,5,000$. This is because the student population of 2019 is smaller than the one in 2014.
The results are broadly in line with the results from 2014. One difference is that we find 2 instead of 3 clusters. This can be explained by the fact that there are much less HBO and WO students present in this student population.
Again, there are stark differences in the bias metric between clusters. For instance, Figure 11(a) shows that in cluster 2, 46% of the students are deemed "high risk", whereas only 17% are deemed "high risk" in cluster 1. This difference is statistically significant at the 0.1% level, as indicated in Table 3 in Appendix C.
@@
Cluster 1 (the cluster with a relatively lower bias) consists exclusively of students following higher professional education (HBO), students that live relatively further away from their parents, and which are relatively older, i.e., exclusively consisting of students older than 21. Cluster 2 (the cluster with a relatively higher bias) consists of 96% of students following vocational education (MBO 1-2, MBO 3-4), students that live relatively closer to their parents, and are relatively younger.
Based on the aggregated data available on the origin of students, as provided by Statistict Netherlands ($n=50,230$) [^16] The aggregated statistics are rounded to the nearest ten; hence, why there are $n=50,230$ reported students instead of ($n=50,233$).. We again observe how particular characteristics used in the risk profile serve as proxies for non-European migration background. Similar to 2014, students following vocational education (MBO 1-2, MBO 3-4) are relatively more likely to have a non-European migration background, and students living relatively closer to their parents, specifically less than 5km, are also relatively more likely to have a non-European migration background. This corresponds to the characteristics of students in cluster 2. Particular combinations of characteristics also observe as proxies for non-European migration background. For instance, similar to 2014, students older then 23 years who are enrolled in vocational education (MBO 1-2, MBO 3-4) are more likely to have a non-European migration background.
@@
Figure 13: The percentage of students with a non-European migration background for the three used profilgin characteristics: (a) type of education, (b) age and (c) distance to parents for students in the CUB 2019 dataset ($n=50,230$), excluding students where the distance is unknown.
@@
Appendix C Details on Statistical Tests
Here, we provide details for the statistical tests used to investigate whether or not differences between the bias metric per cluster are statistically significant. In Table 2 we report results for the student population in 2014, and in Table 3 for the student population in 2019.
@@ Appendix C Details on Statistical Tests
3 6,914 49.46 13.99 35.47 $<10^{-16}$
Table 2: Results of testing difference in bias metrics across clusters returned by the unsupervised bias detection tool on the validation set for the year 2014 ($n=42,920$), using a Pearson’s $\chi^{2}$-test. The ‘$<10^{-16}$’ is used to indicate that the $p-$value was smaller than the smallest decimal that can be represented due to integer underflow.
@@
Table 3: Results of testing difference in bias metrics across clusters returned by the unsupervised bias detection tool on the validation set for the year 2019 ($n=7,796$), using a Pearson’s $\chi^{2}$-test. The ‘$<10^{-16}$’ is used to indicate that the $p-$value was smaller than the smallest decimal that can be represented due to integer underflow. In this case, we only conduct a single statistical test, since there are only two clusters returned.
Appendix D Additional results for Simulation Study
In this section, we report an additional result of the simulation study, complementing the results in Section 5.
In Figure 15 we report the results of using the permutation test when accuracy is used as a bias metric. Unlike the results for using predicted value $\hat{y}_{i}$ as a bias metric in (see Figure 7(b)), we are not very likely to conclude there is a statistically significant difference in the bias metric. Figure 15 also suggests that using the permutation test in this scenario results in the correct probability of falsely rejecting the null hypothesis, given the significance level (5%).
@@ Appendix D Additional results for Simulation Study
Figure 15: The percentage of simulations where for at least one of the clusters $\bar{M}(U_{k})$ is statistically significant from $\bar{M}(U\backslash U_{k})$ at the 5% level for 1,000 simulations for the synthetic dataset with $K=5,n=1,000$ and $d=2$. The error bar reflects the 95% confidence interval. The red dotted line reflects the significance level used when conducting the statistical test for the differences. We measure the difference between $\bar{M}(U\backslash U_{k})$ and $M(U)$ out-of-sample. In Scenario 1 the expectation of $y_{i}$ is the same per cluster. In Scenario 2 the expectation of $y_{i}$ linearly increases per cluster. As a bias metric, we use the accuracy $p(y_{i}\neq\hat{y}_{i})$ rather than directly simulating $m_{i}$, and we use $n_{\mathrm{perm}}=1,000$ permutations.

### v3

@@ Appendix C Applying the unsupervised bias detection tool to the CUB 2019 dataset
We follow the exact same procedure as outlined in Section 4, only changing the options for the minimum number of samples to $1,000,2,000,3,500,\text{and }5,000$. We make this change because the student population of 2019 is smaller than the one in 2014.
The clustering results from the 2019 data are broadly in line with the results from 2014. One difference is that we find 2 instead of 3 clusters. This can be explained by the fact that there are much less HBO and WO students present in this student population.
Again, there are stark differences in the bias metric between clusters. For instance, Figure 8(a) shows that in cluster 2, 46% of the students are deemed "high risk", whereas only 17% are deemed "high risk" in cluster 1. This difference is statistically significant at the 0.1% level, as indicated in Table 4 in Appendix D.
@@
Cluster 1 (the cluster with relatively lower bias) consists exclusively of students following higher professional education (HBO), students who live relatively further away from their parents, and relatively older students, i.e., exclusively consisting of students older than 21. Cluster 2 (the cluster with a relatively higher bias) consists of 96% of students following vocational education (MBO 1-2, MBO 3-4), students who live relatively closer to their parents, and relatively younger students. We include these characteristics found in each cluster as shown in the clustering results in Figure 9.
Based on the aggregated data available on the origin of students, as provided by Statistics Netherlands (n=50,230n=50,230), [^15] The aggregated statistics are rounded to the nearest ten; hence, why there are $n=50,230$ reported students instead of ($n=50,233$). we again observe how particular characteristics used in the risk profile serve as proxies for non-European migration background (see Figure 10). Similar to 2014, students following vocational education (MBO 1-2, MBO 3-4) are relatively more likely to have a non-European migration background, and students living relatively closer to their parents, specifically less than 5km, are also relatively more likely to have a non-European migration background. This corresponds to the characteristics of students in cluster 2. Particular combinations of characteristics also observe as proxies for non-European migration background (see Figure 11). For instance, similar to 2014, students older then 23 years who are enrolled in vocational education (MBO 1-2, MBO 3-4) are more likely to have a non-European migration background.
@@
Figure 10: The percentage of students with a non-European migration background for the three used profiling characteristics: (a) type of education, (b) age and (c) distance to parents for students in the CUB 2019 dataset ($n=50,230$), excluding students where the distance is unknown.
@@
Appendix D Statistical test details and results
Here, we provide details for the statistical tests used to investigate whether or not differences between the bias metric per cluster are statistically significant. After HBAC clustering, we test the following null versus alternative hypotheses for each of the identified clusters:
$H_{0}:\bar{M}(U\backslash U_{k})=\bar{M}(U_{k}),\hskip 7.11317ptH_{A}:\bar{M}(U\backslash U_{k})\neq\bar{M}(U_{k}).$ (1)
For valid statistical inference, the null and alternative hypotheses must be defined before observing the data. If we conduct our statistical test on data that was also used to identify clusters, this condition is violated. Since the clusters are selected to maximize the difference in the bias metric, we will therefore tend to incorrectly conclude that there is bias when there truly is none. The broader issue of testing data-driven hypotheses—often called post-selection inference—has been studied extensively in the context of regularized regression (Kuchibhotla et al., 2022). A long-standing practice in that literature is to use different partitions of the data for model estimation and inference, referred to as sample splitting (Larson, 1931; Stone, 1974). We apply this technique by fitting the HBAC algorithm on a subset of the data and subsequently inferring the clusters and their bias on a new subset of data. Importantly, this set cannot be used to determine our hyperparameters.
When testing for multiple $K>2$ clusters, an issue is that without adjustments, the likelihood of (incorrectly) detecting a statistically significant difference under the null hypothesis increases. We address this by using a Bonferroni correction (Hochberg and Tamhane, 1987).
In Table 3, we report results for the student population in 2014; in Table 4, we provide results for the student population in 2019.
@@ Appendix D Statistical test details and results
3 6,914 49.46 13.99 35.47 $10^{-16}$
Table 3: Results of testing difference in bias metrics across clusters returned by the unsupervised bias detection tool on the validation set for the year 2014 ($n=42,920$), using a Pearson’s $\chi^{2}$-test. The $10^{-16}$’ is used to indicate that the $p$-value was smaller than the smallest decimal that can be represented due to integer underflow.
@@
Table 4: Results of testing difference in bias metrics across clusters returned by the unsupervised bias detection tool on the validation set for the year 2019 ($n=7,796$), using a Pearson’s $\chi^{2}$-test. The $10^{-16}$’ is used to indicate that the $p$-value was smaller than the smallest decimal that can be represented due to integer underflow. In this case, we only conduct a single statistical test, since there are only two clusters returned.
Appendix E Simulation study to evaluate the clustering algorithm
This section provides simulations to support several of the design choices made when applying the Hierarchical Bias-Aware Clustering (HBAC) algorithm, and develops empirical guidance for how to best use it. Similarly to Misztal-Radecka and Indurkhya (2021), we study the performance of the HBAC algorithm on synthetic data. The data-generation process is as follows. Let $\iota_{d}\in\mathbb{R}^{d}$ be a column vector of ones, and $I_{d}\in\mathbb{R}^{d\times d}$ the identity matrix. We use $\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$ to refer to a (multivariate) Gaussian distribution with mean $\boldsymbol{\mu}\in\mathbb{R}^{d}$ and covariance matrix $\boldsymbol{\Sigma}\in\mathbb{R}^{d\times d}$. The features per user for cluster $k$ are generated as:
$m_{i}\sim\mathcal{N}(\eta_{k},1).$ (3)
The $\eta_{k}$ is defined in line with one of two scenarios:
We fit the HBAC algorithm to this synthetic data, and afterwards test if the difference between $\bar{M}(U\backslash U_{k})$ and $\bar{M}(U_{k})$ is statistically significant at a pre-specified significance level. We highlight three recommendations for this test to have the desired properties for detecting true bias, illustrated via results on the synthetic dataset. We illustrate these recommendations when the synthetic data have $K=5$ clusters, $n=1,000$ observations, and $d=2$ features.
In the simulation study, we identify three separate issues with the original implementation by Misztal-Radecka and Indurkhya (2021) that we argued for in Section 3. Next, we outline these issues and propose a solution.
Testing across multiple clusters requires multiple testing correction: Figure 13(a) shows that without Bonferroni correction the false positive rate is much higher than the expected rate from the significance level of the test (i.e., 5% with $\alpha=0.05$), when there is no difference in bias between clusters. This shows that without Bonferroni correction, we might conclude that our classifier is biased, when it is not. With Bonferroni correction, the false positive rate does not exceed the expected rate (i.e., 5%). However, this comes at a cost: Figure 13(a) shows that when the bias increases per cluster, the Bonferroni correction makes it less likely that we detect differences between clusters (i.e., the true positive rate decreases).
Figure 12: Average of the difference between $\bar{M}(U\backslash U_{k})$ and $\bar{M}(U_{k})$ for 1,000 simulations for the synthetic dataset with $K=5,n=1,000$, and $d=2$. The error bar reflects the 95% confidence interval. In-sample refers to the data used to fit the HBAC algorithm, where out-of-sample refers to the data reserved for statistical testing via sample splitting. In Scenario 1 ($m_{i}$ constant), the expectation of $m_{i}$ is the same per cluster and we expect the HBAC algorithm to not find bias in the identified clusters. This is only the case when measuring the difference out-of-sample. In Scenario 2 ($m_{i}$ increasing), the expectation of $m_{i}$ linearly increases per cluster and we expect the HBAC algorithm to find bias both in-sample and out-of-sample.
(a) The effect of using a Bonferroni correction
(b) The effect of using a permutation test
Different bias metrics require different null-hypotheses: So far, our null hypothesis has been that the difference between $\bar{M}(U\backslash U_{k})$ and $M(U)$ is zero. The validity of this null hypothesis depends on the bias metric used. For this part of the simulation study, we do not directly simulate $m_{i}$, but instead simulate the true attribute $y$ as follows: $y_{i}=\mathrm{Bern}(p_{k})$. In the first scenario, we keep $p_{k}=0.5\ \forall k\in\mathcal{K}$, that is: the probability of $y=1$ is constant for all clusters. In the second scenario, we linearly increase the probability that $y=1$ from 0.1 to 0.9, that is: $p_{k}=0.1+0.8\cdot\frac{k-1}{K-1}$. We then fit a logistic regression to predict $y_{i}$. Here, we only consider one bias metric: the predicted value $\hat{y}_{i}$ (usually used to investigate demographic parity). We consider using accuracy as a bias metric for this experiment in Appendix E. When using $\hat{y}_{i}$ as the bias metric, we observe that even when $y_{i}$ has the same distribution per cluster, there are differences in $\hat{y}_{i}$ per cluster. This is because even though $y_{i}$ is independent from $\boldsymbol{x}_{i}$, $\hat{y}_{i}$ is not independent from $\boldsymbol{x}_{i}$. Thus, if we use the null hypothesis that $\bar{M}(U\backslash U_{k})$ and $\bar{M}(U_{k})$ is zero, we are likely to conclude that our classifier is biased. In Figure 13(b), we illustrate this: if $y_{i}$ is constant per cluster, we incorrectly conclude there are statistically significant differences in $\hat{y}_{i}$. An alternative is to use the permutation test as suggested by Ojala and Garriga (2010). To do so, we permute the order of $y_{i}$ in our data to make it independent of $\boldsymbol{x}_{i}$. By repeating this $n_{\mathrm{perm}}$ times, we obtain samples from our null distribution, the distribution of the bias metric when $\boldsymbol{x}_{i}$ is independent of $y_{i}$. We can then perform a statistical test by examining the probability of observing a value of the bias metric under the null distribution. Figure 13(b) shows the result of using such a permutation test. The $t$-test frequently suggests that there is a statistically significant difference in $\hat{y}_{i}$ when $y_{i}$ is constant. This is not true for the permutation test.
Finally, we emphasize that the appropriate null hypothesis depends on the preference of the user. It could be that a user is not interested in differences in $\hat{y}_{i}$ per cluster that are irreducible, e.g., that appear even if $\boldsymbol{x}_{i}$ is independent of $y_{i}$. In this case, a permutation test is more appropriate than a $t$-test.
@@ Appendix E Simulation study to evaluate the clustering algorithm
Figure 14: The percentage of simulations where for at least one of the clusters $\bar{M}(U_{k})$ is statistically significant from $\bar{M}(U\backslash U_{k})$ at the 5% level for 1,000 simulations for the synthetic dataset with $K=5,n=1,000$ and $d=2$. The error bar reflects the 95% confidence interval. The red dotted line reflects the significance level used when conducting the statistical test for the differences. We measure the difference between $\bar{M}(U\backslash U_{k})$ and $M(U)$ out-of-sample. In Scenario 1, the expectation of $y_{i}$ is the same per cluster. In Scenario 2, the expectation of $y_{i}$ linearly increases per cluster. As a bias metric, we use the accuracy $p(y_{i}\neq\hat{y}_{i})$ rather than directly simulating $m_{i}$, and we use $n_{\mathrm{perm}}=1,000$ permutations.
In Figure 14, we report the results of using the permutation test when accuracy is used as a bias metric. Unlike the results for using predicted value $\hat{y}_{i}$ as a bias metric (see Figure 13(b)), we are not very likely to conclude there is a statistically significant difference in the bias metric. Figure 14 also suggests that using the permutation test in this scenario results in the correct probability of falsely rejecting the null hypothesis, given the significance level of 5%.